OpenStudy (anonymous):

can someone help me understand the different parts of Mean Absolute Deviation???

OpenStudy (anonymous):

\[MAD = \frac{ \Sigma |x - \overline{x}| }{ N }\]

OpenStudy (anonymous):

what does the |x - \(\overline{x}\)| mean? N? \(\overline{x}\)???

OpenStudy (anonymous):

\(\Sigma\)?

OpenStudy (anonymous):

I know what they mean somewhat

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

this may help you a little, im not sure though. it heled me http://www.mathsisfun.com/data/mean-deviation.html

OpenStudy (anonymous):

http://www.glencoe.com/sites/pdfs/impact_math/ls9_c1_mean_absolute_deviation.pdf here this is better

OpenStudy (anonymous):

basically, just find the mean of the data given, then you find the range from each number in the set to the mean.. after that, find the mean of those new numbers

OpenStudy (anonymous):

ah that helps. i am getting an example set and seeing if I got it

OpenStudy (anonymous):

alright , feel free to ask any questions that you have. i can try more to help you understand, if needed that is

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So is the formula \(MAD = \dfrac{ \Sigma |x - \overline{x}| }{ N }\) or \(MAD = \dfrac{ \Sigma |x - \mu | }{ N }\)

OpenStudy (anonymous):

ive usually seen it as the first way

OpenStudy (anonymous):

same

OpenStudy (anonymous):

here ill just give you a problem to see if you can do mean absolute deviation ok? itll be VERY simple

OpenStudy (anonymous):

i am guessing they are the same :)

OpenStudy (anonymous):

they are im sure

OpenStudy (anonymous):

oooh and \(\sigma\) is standard deviation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

used in \(\sigma=\sqrt{\Sigma \dfrac{(x-\overline{x})^2}{N}}\)

OpenStudy (anonymous):



OpenStudy (anonymous):